OpenStudy (anonymous):

Given that P = (5, 9) and Q = (13, 12), find the component form and magnitude of PQ? I found <8,3> but I can't figure out if the second one is sqrt73 or just 73

OpenStudy (amistre64):

square the value: (8^2 , 3^3) add them together 64+9 = 72 and take the square root of it: sqrt(72) and simplify

OpenStudy (amistre64):

72 = 9*8 = 9*4*2 sqrt(9.4.2) = sqrt(9) sqrt(4) sqrt(2) = 3*2 sqrt(2)

OpenStudy (amistre64):

forgot how to add lol

OpenStudy (amistre64):

64+9 = 73 .... so sqrt(730

OpenStudy (amistre64):

sqrt(73) my keyboard hates me

OpenStudy (anonymous):

OH! okay so it's just x62+y^2 basically and then take the square root of what you get? LOL that's okay mine does too

OpenStudy (anonymous):

x^2**

OpenStudy (amistre64):

:) yes

OpenStudy (anonymous):

muchas gracias :D you're a life saver, I suck at vectors.. -__-

OpenStudy (amistre64):

vectors are just right triangles at heart ;) good luck

OpenStudy (anonymous):

They're a pain! xD and thanks I need it...lol!